OpenStudy (anonymous):

Multiply. (3+5sqroot2)(4-sqroot2)

OpenStudy (anonymous):

\[(3 + 5\sqrt{2})(4-\sqrt{2})\]You can multiply these out by using the FOIL method. I'll get you started: \[F - (3)(4) = 12\] \[O - (3)(-\sqrt{2}) = -3\sqrt{2}\]